OpenStudy (anonymous):

the slope of the line perpendicular to y=3/4x-7 is a.3/4 b.-4/3 c.4/3 d.-3/4

OpenStudy (anonymous):

slope = -1/m when given slope is m

OpenStudy (earthcitizen):

m2=-4/3

OpenStudy (anonymous):

Assuming 3/4x = (3/4)x (or 3x/4) and not 3/(4x), then (-4/3)x will be perpendicular.

OpenStudy (anonymous):

(I guess it would have to be (3/4)x given the context.)